### v2

@@ Abstract.
We present a fast, robust algorithm for applying a matrix transfer function of a linear time invariant system (LTI) in time domain. Computing $L$ states of a multiple-input multiple-output (MIMO) LTI appears to require $L$ matrix-vector multiplications. We demonstrate that, for any finite user-selected accuracy, the number of matrix-vector multiplications can be reduced to $\mathcal{O}\left(\log_{2}L\right)$ (within an $\mathcal{O}\left(L\right)$ algorithm). The algorithm uses an approximation of the rational transfer function in the z-domain by a matrix polynomial of degree $12^{N+1}-$ 1, where $N$ is chosen to achieve any user-selected accuracy. Importantly, using a cascade implementation in time domain, applying the transfer function requires only $N+1$ matrix-vector multiplications. We note that LTI systems are used in state space models (SSMs) for modeling long range dependencies where $L$ is large. In applications where the state matrix of LTI system is approximated by a structured matrix, the computational cost is further reduced. We briefly describe several structured approximations of matrices that can be used for such purpose.
@@ 1. Introduction
In [1, 2, 3] authors describe an efficient approach to sequence modeling using state space models (SSMs) to address long-range dependencies. While their approach yields an efficient algorithm, we propose a new fast, robust algorithm for the same purpose.
The setup of SSM uses a linear multiple-input multiple-output (MIMO) time invariant system (LTI),
@@
Given the input sequence $\mathbf{u}_{n}$, a computational goal is to obtain the output sequence $\mathbf{y}_{n}$ using the recurrence (1.2). While the recurrence can be used directly, it is prohibitively expensive if the goal is to account for long range dependencies. Alternatively, solving the recurrence, results in a convolution with a transfer function which maps the sequence $\left\{\mathbf{u}_{n}\right\}_{n=0}^{\infty}$ to the sequence $\left\{\mathbf{y}_{n}\right\}_{n=0}^{\infty}$ directly. Using the z-transform (described below for completeness), the transfer function between these two sequences is obtained as
@@
where $I_{m}$ is the $m\times m$ identity matrix. The question is how to apply efficiently the operator $H$ as a convolution in time domain.
As reported in [1], using an example of the so-called HiPPO matrix, diagonalization of the matrix $\overline{A}$ leads to numerical difficulties (i.e. ill-conditioned eigenvectors). As a way to avoid computational issues, the matrix $\overline{A}$ is modified to be structured, i.e. by changing basis it becomes diagonal-plus-low-rank (see also [4]). Using the Fourier transform as a way to accelerate application of a convolution was suggested in [1] as well as in [5].
In this paper we propose a simple fast robust algorithm of applying the operator $H$ in time domain. Additional improvements using structured matrices can be used within the new algorithm to further reduce the computational cost, see Section 5.
@@ 2. Fast cascade algorithm
In [6, 7] we introduced a novel method of accurate approximation of causal and anti-causal IIR filters by a cascade of FIR filters leading to automatic design of efficient digital filters. Many properties, such as filters with exact linear phase or symmetric filters satisfying the perfect reconstruction condition, can only be obtained using non-causal IIR filters. Our method of approximating causal and anti-causal IIR filters produces FIR filters that, within any user-selected accuracy, inherit properties of IIR filters. The resulting cascades of FIR filters have a straightforward parallel implementation. A simple technical tool underpinning these results leads to a fast algorithm for applying the operator $H$ in time domain. For completeness, we present a detailed derivation.
@@
Effectively, we are approximating a rational matrix function $\left(I_{m}-z^{-1}\overline{A}\right)^{-1}$ by a matrix polynomial. If the magnitude of some eigenvalues of the matrix $A$ is close to $1$, then the degree $12^{N+1}-$ 1 of the matrix polynomial $H_{N}$ can be large. However, the cost of applying the convolution operator as a cascade in time domain using this matrix polynomial requires only $N+1$ matrix-vector multiplications. We have
@@
3. Applying convolution via a cascade
In time domain, given an input sequence $\left\{\mathbf{u}_{\ell}\right\}_{\ell=0}^{L-1},$ the representation of the matrix polynomial in (2.3) yields a fast cascade algorithm for applying convolution to obtain the sequence $\left\{\mathbf{y}_{\ell}\right\}_{\ell=0}^{L-1}$. Computing the sequence $\mathbf{y}_{\ell}$ using (1.2) and setting $\mathbf{x}_{-1}=0$ yields
(3.1) $\mathbf{y}_{\ell}=\sum_{k=0}^{\ell}C\overline{A}^{\ell-k}\overline{B}\mathbf{u}_{k}+D\mathbf{u}_{k}.$
Lemma 1 allows us to reuse intermediate matrix-vector products by the matrix $\overline{A}$ and, as a result, we need to apply only the powers $\overline{A}^{2^{n}}$, $n=0,\dots,N-1$. Note that the powers of $z^{-2^{n}}$ in (2.3) correspond to a shift by $2^{n}$ in time domain.
We start with
@@ 3. Applying convolution via a cascade
and form an $m\times L$ matrix
@@
(3.2) $\begin{cases}\mathbf{v}_{\ell}^{\left(1\right)}=\overline{A}\mathbf{v}_{\ell-1}^{\left(0\right)}+\mathbf{v}_{\ell}^{\left(0\right)}&1\leq\ell\leq L-1\\ \mathbf{v}_{\ell}^{\left(0\right)}&\ell=0\end{cases}$
@@
(3.3) $\begin{cases}\mathbf{v}_{\ell}^{\left(2\right)}=\overline{A}^{2}\mathbf{v}_{\ell-2}^{\left(1\right)}+\mathbf{v}_{\ell}^{\left(1\right)}&2\leq\ell\leq L-1\\ \mathbf{v}_{\ell}^{\left(1\right)}&0\leq\ell\leq 1\end{cases}$
@@
(3.4) $\begin{cases}\mathbf{v}_{\ell}^{\left(3\right)}=\overline{A}^{4}\mathbf{v}_{\ell-4}^{\left(2\right)}+\mathbf{v}_{\ell}^{\left(2\right)}&4\leq\ell\leq L-1\\ \mathbf{v}_{\ell}^{\left(2\right)}&0\leq\ell\leq 3,\end{cases}$
@@
$\begin{cases}\mathbf{v}_{\ell}^{\left(n\right)}=\overline{A}^{2^{n-1}}\mathbf{v}_{\ell-2^{n-1}}^{\left(n-1\right)}+\mathbf{v}_{\ell}^{\left(n-1\right)}&2^{n-1}\leq\ell\leq L-1\\ \mathbf{v}_{\ell}^{\left(n-1\right)}&0\leq\ell\leq 2^{n-1}-1.\end{cases}$
@@
As described, the algorithm requires $\mathcal{O}\left(mL\right)$ memory and $\mathcal{O}\left(m^{2}\log L+mL\right)$ arithmetic operations. Using structured matrices, the cost of arithmetic operations can be further reduced to $\mathcal{O}\left(m\log L+mL\right)$ or $\mathcal{O}\left(m\log m\log L+mL\right)$ depending on the selected representation of matrix $\overline{A}$, see Section 5. Noting that sizes of matrices $\overline{B}$, $C$ and $D$, $m\times p$, $q\times m$ and $q\times p$, are typically much smaller than the $m\times m$ size of $\overline{A}$, matrix-vector multiplications by matrices $\overline{B}$, $C$ and $D$ are significantly less expensive and we did not count them in the complexity estimates. Our complexity estimate is for the case $p=q=1$, a single-input single-output (SISO) model as in [1].
While the computational cost of the standard convolution is proportional to the degree $12^{N+1}-$ 1 of the matrix polynomial $H_{N}$, the application of the cascade requires only $N+1$ matrix-vector multiplications. For example, if $N=14$ then the degree of the matrix polynomial is $327672^{N+1}-1=$ 32767 and convolution requires only 15 matrix-vector multiplications.
@@ 4. An example of HiPPO matrix
where $1\leq n,k\leq m$. This is a so-called HiPPO (high-order polynomial projection operators) matrix [9]. Using it as an example and setting $m=100$, we compute
@@
5. Structured representation of matrices
@@
A class of matrices that admit a sparse structured approximation with complexity of matrix-vector product $\mathcal{O}\left(m\right)$ or $\mathcal{O}\left(m\log m\right)$ is very large and such matrices can potentially be considered in setting up the LTI system. If a MIMO system is used, then structured approximations can be used for matrices $B$ or $\overline{B}$ and $C$ and $D$.
@@ 6. Acknowledgment
The author would like to thank IPAM (UCLA) for the invitation to participate in a long program “Mathematics of Intelligences”, Fall 2024.

### v3

@@ Abstract.
We present an unconditionally stable algorithm for applying matrix transfer function of a linear time invariant system (LTI) in time domain. The state matrix of an LTI system used for modeling long range dependencies in state space models (SSMs) has eigenvalues close to $1$. The standard recursion defining LTI system becomes unstable if the $m\times m$ state matrix has just one eigenvalue with absolute value even slightly greater than 1. This may occur when approximating a state matrix by a structured matrix to reduce the cost of matrix-vector multiplication from $\mathcal{O}\left(m^{2}\right)$ to $\mathcal{O}\left(m\right)$ or $\mathcal{O}\left(m\log m\right).$ We introduce an unconditionally stable algorithm that uses an approximation of the rational transfer function in the z-domain by a matrix polynomial of degree $12^{N+1}-$ 1, where $N$ is chosen to achieve any user-selected accuracy. Using a cascade implementation in time domain, applying such transfer function to compute $L$ states requires no more than $2L$ matrix-vector multiplications (whereas the standard recursion requires $L$ matrix-vector multiplications). However, using unconditionally stable algorithm, it is not necessary to assure that an approximate state matrix has all eigenvalues with absolute values strictly less than 1 i.e., within the desired accuracy, the absolute value of some eigenvalues may possibly exceed $1$. Consequently, this algorithm allows one to use a wider variety of structured approximations to reduce the cost of matrix-vector multiplication and we briefly describe several of them to be used for this purpose.
@@ 1. Introduction
In [1, 2, 3] authors describe an approach to sequence modeling using state space models (SSMs) to address long range dependencies. The setup of SSM uses a linear multiple-input multiple-output (MIMO) time invariant system (LTI),
@@
Modeling long range dependencies leads to a dense matrix $\overline{A}$ with absolute values of eigenvalues close to $1$. As reported in [1], using an example of the so-called HiPPO matrix (see Section 4), diagonalization of the matrix $\overline{A}$ yields ill-conditioned eigenvectors. In [1, 3, 2] the matrix $\overline{A}$ is approximated by a structured matrix, i.e. by changing basis it becomes diagonal-plus-low-rank (see also [4]). Alternatively, using the Fourier transform as a way to accelerate application of a convolution was suggested in [1] as well as in [5]. Importantly, using the recurrence (1.2), it is necessary to assure that structured approximation yields a matrix with magnitude of all eigenvalues less than $1$
Given an input sequence $\left\{\mathbf{u}_{n}\right\}_{n=0}^{\infty}$, the goal is to obtain the output sequence $\left\{\mathbf{y}_{n}\right\}_{n=0}^{\infty}$, which can be accomplished by convolution with a transfer function. Using the z-transform (described below for completeness), the transfer function between these two sequences is obtained as
@@
where $I_{m}$ is the $m\times m$ identity matrix. The question is how to apply efficiently the operator $H$ as a convolution in time domain and avoid issues associated with possible accumulation of error when using the recurrence (1.2) directly.
In this paper we propose a simple unconditionally stable algorithm of applying the operator $H$ in time domain. Additional improvements using structured matrices can be used within the new algorithm to further reduce the computational cost, as briefly described in Section 5.
@@ 2. Fast cascade algorithm
In [6, 7] we introduced a novel method of accurate approximation of causal and anti-causal IIR filters by a cascade of FIR filters leading to automatic design of efficient digital filters. Many properties, such as filters with exact linear phase or symmetric filters satisfying the perfect reconstruction condition, can only be obtained using non-causal IIR filters. Our method of approximating causal and anti-causal IIR filters produces FIR filters that, within any user-selected accuracy, inherit properties of IIR filters. The resulting cascades of FIR filters have a straightforward parallel implementation. A simple technical tool underpinning these results leads to a fast algorithm for applying the operator $H$ in (1.3) in time domain. For completeness, we present a detailed derivation.
@@
Effectively, we are approximating a rational matrix function $\left(I_{m}-z^{-1}\overline{A}\right)^{-1}$ by a matrix polynomial. If the magnitude of some eigenvalues of the matrix $A$ is close to $1$, then the degree $12^{N+1}-$ 1 of the matrix polynomial $H_{N}$ can be large. However, the cost of applying this matrix polynomial as a cascade in time domain requires only $N+1$ matrix-vector multiplications. We have
@@
3. Convolution via a cascade algorithm
Given an input sequence $\left\{\mathbf{u}_{\ell}\right\}_{\ell=0}^{L-1},$ the representation of the matrix polynomial in (2.3) yields a fast cascade algorithm for applying convolution to obtain the sequence $\left\{\mathbf{y}_{\ell}\right\}_{\ell=0}^{L-1}$. Computing the sequence $\mathbf{y}_{\ell}$ using (1.2) and setting $\mathbf{x}_{-1}=0$ yields
(3.1) $\mathbf{y}_{\ell}=\sum_{k=0}^{\ell}C\overline{A}^{\ell-k}\overline{B}\mathbf{u}_{k}+D\mathbf{u}_{k},$
an expression that uses all powers of matrix $\overline{A}$. Lemma 1 allows us to reuse intermediate matrix-vector products by the matrix $\overline{A}$ and, as a result, we need to apply only the powers $\overline{A}^{2^{n}}$, $n=0,\dots,N-1$. Note that the powers of $z^{-2^{n}}$ in (2.3) correspond to a shift by $2^{n}$ in time domain.
We start algorithm by computing
@@ 3. Convolution via a cascade algorithm
and forming an $m\times L$ matrix
@@
(3.2) $\begin{cases}\mathbf{v}_{\ell}^{\left(1\right)}=\overline{A}\mathbf{v}_{\ell-1}^{\left(0\right)}+\mathbf{v}_{\ell}^{\left(0\right)}&1\leq\ell\leq L-1\\ \mathbf{v}_{\ell}^{\left(1\right)}=\mathbf{v}_{\ell}^{\left(0\right)}&\ell=0\end{cases}$
@@
(3.3) $\begin{cases}\mathbf{v}_{\ell}^{\left(2\right)}=\overline{A}^{2}\mathbf{v}_{\ell-2}^{\left(1\right)}+\mathbf{v}_{\ell}^{\left(1\right)}&2\leq\ell\leq L-1\\ \mathbf{\mathbf{v}_{\ell}^{\left(2\right)}=v}_{\ell}^{\left(1\right)}&0\leq\ell\leq 1\end{cases}$
@@
(3.4) $\begin{cases}\mathbf{v}_{\ell}^{\left(3\right)}=\overline{A}^{4}\mathbf{v}_{\ell-4}^{\left(2\right)}+\mathbf{v}_{\ell}^{\left(2\right)}&4\leq\ell\leq L-1\\ \mathbf{\mathbf{v}_{\ell}^{\left(3\right)}=v}_{\ell}^{\left(2\right)}&0\leq\ell\leq 3,\end{cases}$
@@
$\begin{cases}\mathbf{v}_{\ell}^{\left(n\right)}=\overline{A}^{2^{n-1}}\mathbf{v}_{\ell-2^{n-1}}^{\left(n-1\right)}+\mathbf{v}_{\ell}^{\left(n-1\right)}&2^{n-1}\leq\ell\leq L-1\\ \mathbf{v}_{\ell}^{\left(n\right)}=\mathbf{v}_{\ell}^{\left(n-1\right)}&0\leq\ell\leq 2^{n-1}-1.\end{cases}$
@@
As described, the algorithm requires $\mathcal{O}\left(mL\right)$ memory and $\mathcal{O}\left(m^{2}\log L+mL\right)$ arithmetic operations. Using approximation of matrix $\overline{A}$ by a structured matrix, the cost of arithmetic operations can be further reduced to $\mathcal{O}\left(m\log L+mL\right)$ or $\mathcal{O}\left(m\log m\log L+mL\right)$ depending on the selected representation of matrix $\overline{A}$, see Section 5. Noting that sizes of matrices $\overline{B}$, $C$ and $D$, $m\times p$, $q\times m$ and $q\times p$, are typically much smaller than the $m\times m$ size of $\overline{A}$, matrix-vector multiplications by matrices $\overline{B}$, $C$ and $D$ are significantly less expensive and we did not count them in the complexity estimates. Our complexity estimate is for the case $p=q=1$, a single-input single-output (SISO) model as in [1].
While the degree of matrix polynomial $H_{N}$ in (2.3) is $12^{N+1}-$ 1 , the cascade algorithm requires only $N+1$ matrix-vector multiplications. For example, if $N=14$ then the degree of the matrix polynomial is $327672^{N+1}-1=$ 32767, whereas applying it as a convolution requires only 15 matrix-vector multiplications. While providing an approximation with any finite user-selected accuracy, the cascade algorithm is unconditionally stable in contrast with that using recurrence in (1.2) directly.
@@ 4. An example of HiPPO matrix
where $1\leq n,k\leq m$. This is a so-called HiPPO (High-order Polynomial Projection Operator) matrix, see [9]. Using it as an example and setting $m=100$, we compute
@@
5. Structured approximations of matrices
@@
A class of matrices that admit a sparse structured approximation with complexity of matrix-vector product $\mathcal{O}\left(m\right)$ or $\mathcal{O}\left(m\log m\right)$ is large and such matrices can potentially be considered in setting up the LTI system. If a MIMO system is used, then structured approximations can also be used for matrices $B$ or $\overline{B}$ and $C$ and $D$.
6. Discussion
The algorithm described in this paper is formally slower than the direct use of the recurrence (1.2) by a factor not exceeding two ($\overline{A}$ is applied to every entry of the sequence, $\overline{A}^{2}$ to every second entry, $\overline{A}^{4}$ to every forth entry, etc.). This raises a question “what are the advantages of replacing the recurrence (1.2) with a cascade convolution in Section 3?”. If the absolute values of all eigenvalues of the matrix $\overline{A}$ in (1.2) are significantly smaller than $1$, then it is likely that the eigenvalues of a structured matrix constructed as its approximation will also be smaller than $1$. In such cases the recurrence (1.2) is stable and it is not necessary to use the algorithm in Section 3. However, a structured approximation of a matrix $\overline{A}$ with absolute values of eigenvalues close to $1$ may end up with eigenvalues that violate this condition (within the accuracy of approximation). In such cases the recurrence in (1.2) is unstable whereas the cascade convolution in Section 3 is always stable. Note that in applications with a goal of analyzing long range dependencies, at least some eigenvalues of a matrix $\overline{A}$ will have absolute value close to $1$. In such problems, with a modest increase in computational cost, the algorithm in Section 3 provides for a wider choice of possible structured approximations and allows one to consider alternative bases than those used in [1, 8, 2, 4].
@@ 7. Acknowledgment
Part of this research was performed while the author was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919).